OpenStudy (anonymous):

critical thingking three compound containing potassium and oxygen are compared. analysis shows that for each 1.00g of 0, the compounds have 1,22g, 2.44g, and 4.89g of potassium respectively. show how these data support the law of multiple proportions. what is the answer?

OpenStudy (frostbite):

I can just give you the answer, but I'll like to help get to it. I can do the first one, then you can do the two others. We see there is a relationship between oxygen and potassium which respect to content. that mean there most also be a relationship between their amount of different atoms. We use the relation \[\large n=\frac{ m }{ M }\]To find out how many moles we got of the compound. We can start with potassium: \[\large n(K)=\frac{ m(K) }{ M(K) }=\frac{ 1.22 ~ gram }{ 39.1 ~ \frac{ gram }{ mol } }=0.0312 ~ mol\] The ratio will then become: \[\large ratio = \frac{ 0.0312 ~ mol }{ 0.0312 ~ mol }=1\] For oxygen we do the same: \[\large n(O)=\frac{ m(O) }{ M(O) }=\frac{ 1 ~ g }{ 16.0 \frac{ gram }{ mol } }=0.0625 ~ mol\] Ratio: \[Ratio=\frac{ 0.0625 ~ mol }{ 0.0312 ~ mol }\approx 2\] So the compound most be \(KO_{2}\), which confirm law of multiple proportions.

OpenStudy (frostbite):

Try do the 2 others, if you need help I'll be here.